Give the Relation between refractive index of a medium and wavelength

Grade:12

2 Answers

Umakant biswal
7 years ago
dear mohana 
v = lambda * frequency 
so, by combining we get 
lambda(0) / lambda = c/f / v/f = f*c / f*v  = c/v = n 
hope it clears now 
regards 
Soham Borawake
7 years ago
Refractive Index for eg of glass wrt. Water is wUg. And wavelength is lamda(water) and lamda(glass)... There fore wUg= lamda (glass)/lamda (water)